Allen is 25 years older than Beth. In 10 years, Allen will be twice as old as Beth. How old are Allen and Beth right now?

Answer by greenestamps(13198)



Algebraically....

x = Beth's age
x+25 = Allen's age

x+10 = Beth's age in 10 years
x+35 = Allen's age in 10 years

x+35 = 2(x+10)

Solve using basic algebra.

By logical reasoning (good mental exercise, if an algebraic solution is not required)....

The difference in their ages will always be 25 years.
So when Allen is twice as old as Beth, their ages will be 25 and 50.
Since that will be 10 years from now, their current ages are 15 and 40.